OpenStudy (anonymous):

If a circle has the diameter of 8, what is the circumference? 16.28 50.24 18.41 25.13

OpenStudy (mitu12):

25.13

OpenStudy (jdoe0001):

\(\bf circumference=2\pi {\color{brown}{ r}}\qquad {\color{brown}{ r}}=radius=\frac{diameter}{2}\)

OpenStudy (igreen):

@jdoe0001 Or \(c = \pi d\)

OpenStudy (igreen):

Where d = diameter.

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

thank you guys!!!!!!!!!

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @JRDetroit